OpenStudy (anonymous):

can someone solve this ? |x+1|^2 + 2 |x+2| ≥ 2

OpenStudy (michele_laino):

we can note this: \[\Large {\left| {x + 1} \right|^2} = {\left( {x + 1} \right)^2}\]

OpenStudy (empty):

There's really no difference between \(|x+1|^2\) and \((x+1)^2\) since they both evaluate to the same thing, so you can safely throw these absolute value signs away

OpenStudy (empty):

welp

OpenStudy (michele_laino):

so, we can rewrite your inequality, as below: \[\Large {\left( {x + 1} \right)^2} + 2\left| {x + 2} \right| \geqslant 2\]

OpenStudy (anonymous):

so @Michele_Laino (x+1)^2 = x^2+2x+1 ?

OpenStudy (michele_laino):

next, we have to consider the subsequent two cases: \[\large \left\{ \begin{gathered} {\left( {x + 1} \right)^2} + 2\left( {x + 2} \right) \geqslant 2 \hfill \\ x + 2 \geqslant 0 \hfill \\ \end{gathered} \right.,\quad \left\{ \begin{gathered} {\left( {x + 1} \right)^2} - 2\left( {x + 2} \right) \geqslant 2 \hfill \\ x + 2 < 0 \hfill \\ \end{gathered} \right.\] please solve both systems above

OpenStudy (madhu.mukherjee.946):

did you solve it